OpenStudy (anonymous):

Find the zeros of the function: 8c^2=4c

OpenStudy (mertsj):

1. Subtract 4c from both sides. 2. Factor 3. Set each factor equal to 0 4. Solve each equation.

OpenStudy (anonymous):

im sorry im still confused

OpenStudy (anonymous):

@ineedyouubiebs what he meant to say was 8c^2 - 4c = 4c - 4c so 8c^2 - 4c = 0 Factorise this expression.

OpenStudy (anonymous):

@luckythebest so would it be 4c(2c-c) ????????

OpenStudy (anonymous):

@ineedyouubiebs not really It will be 8c^2-4c = 0 4c(2c - 1) = 0 , and not 2c-c

OpenStudy (anonymous):

and my next step would be what?

OpenStudy (mertsj):

Set each factor equal to 0

OpenStudy (anonymous):

@luckythebest no what you said was true

OpenStudy (anonymous):

@ineedyouubiebs if 4c(2c-1) = 0, then one of these two will be 0. 4c = 0 or 2c-1 = 0.

OpenStudy (anonymous):

so then my final answer would just be 0 and 0?

OpenStudy (anonymous):

@ineedyouubiebs nope. we have two equations whose solution is 0. so 4c = 0 and 2c-1 = 0 i.e. c = 0/4 or 2c = 1. accordingly solve for c. :)